OpenStudy (anonymous):

The enrollment at a local college increased 8% over last year's enrollment of 900. Find the current enrollment.

OpenStudy (anonymous):

900 *1.08=972 enrollments.

OpenStudy (anonymous):

10%=90 1%=9 9x8=72 72+900=972

OpenStudy (anonymous):

current = 900 + 900* 8% = 972 students